# SpatialLM: Training Large Language Models for Structured Indoor Modeling

### v1

@@ C.1 Layout Estimation
RoomFormer [60] 70.4 86.0 69.9 55.2 67.2 84.6 66.4 50.7
SceneScript [2] 83.1 92.9 81.6 74.7 80.8 91.9 79.7 70.9
SpatialLM (ft. Structured3D) 32.8 54.1 25.0 19.4 17.9 33.5 12.0 8.3
SpatialLM (ft. Ours) 51.2 71.5 33.3 48.7 38.3 64.8 29.0 21.0
SpatialLM (ft. Ours $\rightarrow$ Structured3D) 86.5 95.3 83.7 80.4 84.6 94.5 81.5 77.7
@@
V-DETR [47] 65.1 77.4 80.1 48.8 44.4 82.5 55.6 66.1 68.0 62.3 54.0 47.4 50.8 74.0 79.5 72.7 55.3 98.0 54.9
SceneScript [2] 49.1 64.5 71.1 39.7 30.3 81.1 43.4 30.9 53.4 41.7 42.6 11.8 28.2 58.6 48.9 68.3 55.8 77.5 36.5
@@
V-DETR [47] 56.8 68.8 75.5 50.4 39.1 76.2 43.6 54.1 52.1 50.9 47.8 42.7 49.2 61.5 61.4 65.4 50.6 88.6 43.9
SceneScript [2] 36.8 54.8 65.9 37.3 21.2 73.7 25.3 16.9 44.5 23.5 30.5 3.0 23.3 19.5 24.9 62.0 50.1 64.2 21.7
@@ C.3 Zero-shot Detection on Videos
Per-category results. We collect 107 virtual room-tour videos and use MASt3R-SLAM [36] to reconstruct a point cloud from each video. Note that these videos are rendered from virtual scenes created by professional designers, thus ground-truth 3D annotations are available. To calibrate the reconstructed point cloud with ground truth 3D annotations, we align the estimated camera trajectory from MASt3R-SLAM to the ground-truth camera trajectory. Then, we perform zero-shot detection with SpatialLM on the calibrated point clouds.
In Table 12, we report the quantitative results on the top-20 occurring categories in the video test set. Note that this is a challenging task, because the reconstructed point clouds are often noisy and highly incomplete. Furthermore, because of the differences in scale and position due to imperfect alignment of camera trajectories, the objects in the point cloud may not match well with the ground truth 3D annotations, leading to lower F1 scores.
@@ D.1 Supporting Point Clouds from Diverse Sources
  • Text-to-3D: We first generate an image via Flux-dev [26] with prompt “a low poly 3D living room in cartoon style”, then convert the image into 3D using TRELLIS [57].
  • Hand-held video camera: We take a real-world video using a hand-held camera and perform 3D reconstruction using MASt3R-SLAM [36].
  • LiDAR-based reconstruction: We take point clouds reconstructed from captures using iPhone ARKit from MultiScan [34].
  • Synthetic mesh: We create point clouds by randomly sampling mesh surfaces from a synthetic dataset HSSD [24].
As can be seen in Figure 12, our model is robust to variations in point cloud origin, structure, and appearance.
@@ D.2 Task Adaptation via Language-based Prompts
Detection with user-specified categories. We enable the model to perform object detection conditioned on user-specified categories by leveraging the flexibility of LLMs. During training, the input prompt is customized to indicate which object categories to detect. The model then learns to selectively predict oriented bounding boxes (OBBs) for instances matching the specified categories, all within a unified training procedure. This experiment highlights the model’s adaptability through prompt-based control. Some qualitative results are provided in Figure 13.
@@
Compared to 3D object detection, this task is relatively easy. Nevertheless, it demonstrates the versatility and potential use of large language models for structured 3D understanding. We show some qualitative results in Figure 14. Our model achieves an overall classification accuracy of 96.8% on the test set of our dataset.

### v2

@@ C.1 Layout Estimation
RoomFormer [YueKSE23] 83.4 86.0 85.1 78.9 81.4 84.6 83.1 76.6
SceneScript [SceneScript] 90.4 92.9 92.8 85.5 89.2 91.9 92.5 83.3
SpatialLM (ft. Structured3D) 32.6 54.1 19.4 24.3 18.1 33.5 9.4 11.5
SpatialLM (ft. Ours) 59.9 71.5 42.7 65.6 44.7 64.8 39.7 29.6
SpatialLM (ft. Ours $\rightarrow$ Structured3D) 94.3 95.3 95.9 91.7 93.5 94.5 95.7 90.2
@@
V-DETR [0001GYLLW00G24] 65.1 77.4 80.1 48.8 44.4 82.5 55.6 66.1 68.0 62.3 54.0 47.4 50.8 74.0 79.5 72.7 55.3 98.0 54.9
SceneScript [SceneScript] 49.1 64.5 71.1 39.7 30.3 81.1 43.4 30.9 53.4 41.7 42.6 11.8 28.2 58.6 48.9 68.3 55.8 77.5 36.5
@@
V-DETR [0001GYLLW00G24] 56.8 68.8 75.5 50.4 39.1 76.2 43.6 54.1 52.1 50.9 47.8 42.7 49.2 61.5 61.4 65.4 50.6 88.6 43.9
SceneScript [SceneScript] 36.8 54.8 65.9 37.3 21.2 73.7 25.3 16.9 44.5 23.5 30.5 3.0 23.3 19.5 24.9 62.0 50.1 64.2 21.7
@@ C.3 Zero-shot Detection on Videos
Per-category results. We collect 107 virtual room-tour videos and use MASt3R-SLAM [MAST3RSLAM] to reconstruct a point cloud from each video. Note that these videos are rendered from virtual scenes created by professional designers, thus ground-truth 3D annotations are available. To calibrate the reconstructed point cloud with ground truth 3D annotations, we align the estimated camera trajectory from MASt3R-SLAM to the ground-truth camera trajectory. Then, we perform zero-shot detection with SpatialLM on the calibrated point clouds.
In Table˜12, we report the quantitative results on the top-20 occurring categories in the video test set. Note that this is a challenging task, because the reconstructed point clouds are often noisy and highly incomplete. Furthermore, because of the differences in scale and position due to imperfect alignment of camera trajectories, the objects in the point cloud may not match well with the ground truth 3D annotations, leading to lower F1 scores.
@@ D.1 Supporting Point Clouds from Diverse Sources
  • Text-to-3D: We first generate an image via Flux-dev [flux2024] with prompt “a low poly 3D living room in cartoon style”, then convert the image into 3D using TRELLIS [trellis].
  • Hand-held video camera: We take a real-world video using a hand-held camera and perform 3D reconstruction using MASt3R-SLAM [MAST3RSLAM].
  • LiDAR-based reconstruction: We take point clouds reconstructed from captures using iPhone ARKit from MultiScan [MaoZJCS22].
  • Synthetic mesh: We create point clouds by randomly sampling mesh surfaces from a synthetic dataset HSSD [HSSD].
As can be seen in Figure˜13, our model is robust to variations in point cloud origin, structure, and appearance.
@@ D.2 Task Adaptation via Language-based Prompts
Detection with user-specified categories. We enable the model to perform object detection conditioned on user-specified categories by leveraging the flexibility of LLMs. During training, the input prompt is customized to indicate which object categories to detect. The model then learns to selectively predict oriented bounding boxes (OBBs) for instances matching the specified categories, all within a unified training procedure. This experiment highlights the model’s adaptability through prompt-based control. Some qualitative results are provided in Figure˜14.
@@
Compared to 3D object detection, this task is relatively easy. Nevertheless, it demonstrates the versatility and potential use of large language models for structured 3D understanding. We show some qualitative results in Figure˜15. Our model achieves an overall classification accuracy of 96.8% on the test set of our dataset.